OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (anonymous):

First rearrange the top equation so you have x in terms of y and then sub this into the second equation.

OpenStudy (anonymous):

so slope intercept form?

OpenStudy (anonymous):

y=mx+b ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

7x-2y=21 7x 7x 2y= 7x+21?

OpenStudy (anonymous):

2y = 7x -21

OpenStudy (anonymous):

Okay , then plug it into the 2nd equation??

OpenStudy (anonymous):

yeah, then you will have an equation in terms of x only and can solve for x. Then take this value of x and plug back into the first equation to solve for y.

OpenStudy (anonymous):

right right right , okay . So solving it would it be like this? 2y= 7x- 21 2 2 2 y= 7/2x - 10.5 ? but that would give it decimals .. and none of the options have a decimal , where am i messing up at .

OpenStudy (anonymous):

I suggest you add the equations, then isolate why and find its value. After doing so, substitute it in y and you will get x. Hope I helped.

OpenStudy (anonymous):

You could do it the other way round and solve the bottom one for y = 57 - 4x then sub this into the top one that way there will be no decimals. Either way you should end up with whole numbers by the end

OpenStudy (anonymous):

so confused.

OpenStudy (anonymous):

Ive tried literally every way you guys have told me and not gotten none of those 4 answers... /:

OpenStudy (anonymous):

If you rearrange the bottom equation to get y=57 -4x then you can sub this back into the top one to get 7x - 2(57-4x) = 21

OpenStudy (anonymous):

Okay i got 7x-114-8x=21

OpenStudy (anonymous):

I dont know what to do from here .

OpenStudy (anonymous):

Dont know how to do that.

OpenStudy (anonymous):

You need to get all the x terms on one side and all the numbers on the other. Add 114 to both sides then do 7x-8x for the left hand side.

OpenStudy (anonymous):

Im just gonna guess the answer.Thanks for trying to help tho

OpenStudy (anonymous):

lol fair enough

OpenStudy (anonymous):

x=9 y = 21